OpenStudy (shyannep913):

please help Yulian works at the zoo feeding the animals. He puts water in the elephant habitat at the beginning of the day. This table shows how much water, in gallons, remains at different times during the day. Hours Water Remaining 2 43 4 36 6 29 8 22 What is the equation that represents Yulian’s situation?

OpenStudy (excalibur0126):

Every two hours, the elephants drink 7 gallons of water. So, you can tell that he gives them 50 gallons of water at the start of the day, the y intercept. The slope you can figure out with the handy equation (y2-y1)/(x2-x1). For example, we'll be using (2,43) and (4,36), with the second set of coordinates being (x2, y2). 36-43=-7, 4-2=2. -7/2=-3.5, the slope. The slope is what x is multiplied by and the y intercept is added after x. So the final equation would be y=-3.5x+50.